OpenStudy (anonymous):

Solve the following system of equations: -2x + y = 1 -4x + y = -1 (3, 1) (-1, 3) (-1, -3) (1, 3)

OpenStudy (radar):

Since there is no direction as to what method, this system looks good to use the "elimination" method. In using the elimination, I like to add to eliminate. but this problem is easier to subtract. Subtract equation no. 2 from no. 1 -2x + y = 1 -(-4x + y =-1) Subtracting --------------- 2x = 2 x = 1 Now substitute the value of x (1) in the first equation and solve for y. -2 + y = 1 y=1 + 2=3 y=3 verify (it checks)